OpenStudy (anonymous):

find the missing value, the principal is 15,300 the rate is 3.5% the simple interest is 2,142

OpenStudy (anonymous):

you have to find out the number of years

OpenStudy (anonymous):

\[Amount = Principal\times{Interest rate}\times{Time}\]

OpenStudy (anonymous):

amount = Principal x (1+(interest rate x time))

OpenStudy (anonymous):

I=PRT 3.5% = 3.5/100 = 0.035 therefore, 2142 = 15300 x 0.035 x T 2142 = 535.5 x T T = 2142/535.5 T= 4 years

OpenStudy (anonymous):

does that make sense? @triciahickman

OpenStudy (anonymous):

thamk you

OpenStudy (anonymous):

no problem! c: